OpenStudy (anonymous):

Evaluate the function f(x)= f(x)=2x-2, when x=-1

OpenStudy (nurali):

f(x)=2x-2, when x=-1 f(-1)=2(-1)-2 f(-1)=-2-2 f(-1)=-4

OpenStudy (anonymous):

So what you want to do is plug -1 in for x and evaluate, then the answer you get from the first equation you plug in again since it says f(x)=f(x).